# Enhancing Consistency and Mitigating Bias: A Data Replay Approach for Incremental Learning

### v1

@@ IV-B1 Estimate Data Consistency
TABLE I: Effects of Different Losses on Data Consistency. We ablate loss components of the baseline loss and show the improvement after combining $L_{dce}$.
@@
To estimate $p(\boldsymbol{z})$ and $q(\boldsymbol{z})$, we adopt the kernel density estimation (KDE) technique. Specifically, we use the Gaussian kernel and decide the bandwidth by Scott’s method [50]. With estimated $p(\boldsymbol{z})$ and $q(\boldsymbol{z})$, one can calculate $D_{KL}$. But owing to no analytical solution and high computational complexity, we approximate KL divergence with Monte-Carlo approximation [^1] We follow the algorithm in http://joschu.net/blog/kl-approx.html.. Though performing estimation in this way is feasible and the estimated KL divergence can be used as a metric, it is inefficient in computing and is not differential to the inverted samples in the inversion stage which incurs difficulties when adopting it as the objective function.
To fulfill the vision of efficiency and differentiability, we further assume that the class-conditional distribution in the feature space follows a tied multivariate Gaussian distribution. It is reasonable for the theoretical connection between Gaussian discriminant analysis and the SoftMax classifier and the empirical results detailed in  [43]. To be specific, given any class $k$, denote the feature set of class $k$ as $Z_{k}$ satisfying: $Z_{k}\sim\mathcal{N}(\boldsymbol{u_{k}},\boldsymbol{\bar{\Sigma}})$, where $\boldsymbol{u_{k}}$ is the mean vector and $\boldsymbol{\bar{\Sigma}}$ is the tied covariance matrix. The parameters can be estimated easily: $\boldsymbol{u_{k}}=\frac{1}{|Z_{k}|}\sum_{\boldsymbol{z}\in Z_{k}}\boldsymbol{z}$, $\boldsymbol{\bar{\Sigma}}=\frac{1}{|Z|}\left(\sum_{k}{\sum_{\boldsymbol{z}\in Z_{k}}\left(\boldsymbol{z}-\boldsymbol{u_{k}}\right)\left(\boldsymbol{z}-\boldsymbol{u_{k}}\right)^{T}}\right)$, where $Z=\bigcup\limits_{k}Z_{k}$. In addition, $p(k)$ can be estimated by $|Z_{k}|/|Z|$. $p(\boldsymbol{z})$ is then given as follows:
$q(\boldsymbol{z})$ can be estimated similarly is spite that $\hat{Z}$ is the feature set of synthetic samples. Because there is no analytical form of $D_{KL}$ between the mixtures of Gaussians, we adopt the Monte-Carlo approximation method as above.
IV-B2 Analysis of Existing Losses
With the two KL Divergence metrics, we try to study the effects of losses proposed in the literature under a unified framework. We first train a model with 50 classes in CIFAR-100 and then invert samples from it. Then, the KL divergences of inverted and real old data approximated in two ways are given. We ablate the loss components of the inversion baseline and the results can be found in Table I.
@@ IV-B2 Analysis of Existing Losses
We can draw some conclusions from Table I. First, the results of $D_{KL}$ (KDE) and $D_{KL}$ (Gaussian) are similar, which corroborate each other and can reflect the relative data consistency among methods. Second, the cross-entropy loss does not contribute to data consistency much. It is intuitive because cross-entropy loss is a sample-wise measurement and lacks any direct constraint on either the diversity or the distribution. What is more, minimizing the cross-entropy loss in Equation 1 equals maximizing the Softmax score of samples. But it is well-studied that Softmax score [51] is not a good metric for OOD detection [45, 44] and thus for data consistency measurement. Note in our experiments, the cross-entropy is not indispensable. We keep it for it does not contradict other losses during training and can slightly improve the performance in some cases. Third, the statistics alignment loss indeed ameliorates the consistency, but there is still plenty of room. We observe that the alignment in Equation 2 only considers the features of the BN layers but lacks explicit regularization on non-BN layer features, such as features of the important penultimate layer. To that effect, we propose the data consistency enhancement loss $L_{dce}$ on the features of the penultimate layer described below. Combining $L_{dce}$, the data consistency is further enhanced as desired.
IV-B3 Data Consistency Enhancement Loss
In what follows, we give the details of $L_{dce}$, which intends to align the distributions of real and synthetic data under the aforementioned Multivariate Gaussian assumption. Formally, given any class $k\in\mathcal{C}_{0:N-1}$ and its data $X_{k}$, denote the feature vectors of $X_{k}$ as $h(X_{k})$ and $h(X_{k})$ satisfies: $h(X_{k})\sim\mathcal{N}(\boldsymbol{u_{k}},\boldsymbol{\bar{\Sigma}})$. In the incremental learning setting, we estimate $\boldsymbol{u_{0}},\boldsymbol{u_{1}},\dots,\boldsymbol{u_{|\mathcal{C}_{0:i}|}}$ and $\boldsymbol{\bar{\Sigma}}$ at the end of each task $\mathcal{T}_{i}$ with $h_{0:i}(\cdot)$ as follows:
$\boldsymbol{u_{k}}=\begin{cases}\frac{1}{|\hat{X}_{k}^{i-1}|}\sum\limits_{\hat{x}\in\hat{X}_{k}^{i-1}}h_{0:i}(\hat{x}),&\text{if $k\notin\mathcal{C}_{i}$,}\\ \frac{1}{|X_{k}|}\sum\limits_{x\in X_{k}}h_{0:i}(x),&\text{if $k\in\mathcal{C}_{i}$,}\end{cases}$ (8)
$\boldsymbol{\bar{\Sigma}}=$ $\frac{1}{|\hat{X}^{i-1}\cup X^{i}|}\Big{(}\sum_{k\notin\mathcal{C}_{i}}{\sum_{\hat{x}\in\hat{X}_{k}^{i-1}}\left(h_{0:i}(\hat{x})-\boldsymbol{u_{k}}\right)\left(h_{0:i}(\hat{x})-\boldsymbol{u_{k}}\right)^{T}}$
$+\sum_{k\in\mathcal{C}_{i}}{\sum_{x\in X_{k}}\left(h_{0:i}(x)-\boldsymbol{u_{k}}\right)\left(h_{0:i}(x)-\boldsymbol{u_{k}}\right)^{T}}\Big{)},$ (9)
where $\hat{X}^{i-1}$ denotes the data inverted from $f_{0:i-1}$ and $\hat{X}_{k}^{i-1}$ denotes the inverted data of class $k$. Then, to align the synthetic data $\hat{X}^{i}$ in task $\mathcal{T}_{i+1}$ and data $\hat{X}^{i-1}\cup X^{i}$ in task $\mathcal{T}_{i}$, we estimate corresponding parameters of $\hat{X}^{i}$ as follows:
$\boldsymbol{\hat{u}_{k}}=\frac{1}{|B_{k}|}\sum\limits_{\hat{x}\in\hat{X}_{k}^{i}\cap B}h_{0:i}(\hat{x})$ (10)
$\boldsymbol{\hat{\Sigma}}=\frac{1}{|B|}\sum_{k}{\sum_{\hat{x}\in\hat{X}_{k}^{i}\cap B}\left(h_{0:i}(\hat{x})-\boldsymbol{\hat{u}_{k}}\right)\left(h_{0:i}(\hat{x})-\boldsymbol{\hat{u}_{k}}\right)^{T}},$ (11)
Figure 3: The norms of class weights in the standard 5-task setting.
IV-C Debiased Classifier
@@
Figure 4: The norms of class weights in 2-task experiments with different training schemes for the second task. (a) the standard training scheme. (b) a scheme where real old data is used and trained with local CE loss. (c) a scheme where the real old data is used. (d) a scheme where the synthetic old data is trained with the local CE loss. We report the results at the end of the final task.
By treating real and synthetic samples differently, the works of [12, 13] successfully circumvented the domain classification problem and achieved high performance. However, based on our observation, it incurs a class weight bias between old and new classes: the norm of weights of old classes is smaller than that of new classes and the gap continually widens as learning progresses. Denote the parameters in $g(\cdot)$ as $\boldsymbol{W}\in\mathbb{R}^{K\times d}$, where $d$ is the dimension of feature $h(x)$ and $K$ is the number of classes. $k$-th row vector in $\boldsymbol{W}$ is denoted as $\boldsymbol{w_{k}}$, which is the weight vector corresponding to class $k$. We plot the weight norm versus the class order of tasks in Figure 3. We analyze the underlying reasons and propose a simple solution to the problem below.
IV-C1 Analysis of Weight Bias
@@
However, when we control the number of classes of synthetic and real data to be the same, the weight bias problem still exists (as shown in Figure 4(a)). It indicates the existence of other reasons. We argue different loss functions and different training data may also contribute to weight bias and thus design a two-task class incremental experiment to uncover the factors concerned with it. The first task includes data from 50 classes of CIFAR-100 while the second task includes the other 50 classes. Training with standard CE is adopted for the first task. As for the second task, we compare 4 training schemes: (1) synthetic (HKD) - real (LCE), which is the standard training scheme, (2) real (LCE) - real (LCE), where real old data is used and trained with local CE loss, (3) real (HKD) - real (LCE), where the real old data is used, (4) synthetic (LCE) - real (LCE), where the synthetic old data is trained with the local CE loss. We report the norm of class weights at the end of task two.
First of all, Figure 4(b) shows an unbiased situation under the condition of equal losses and data quality for both tasks. Further, results in Figure 4(c) and 3(d) validate our hypothesis that both the different training losses and different properties of training data will have effects on the class weights. HKD loss calculates the mean square error (MAE) of the logits of the old and new models as defined in Equation 5. LCE loss defined in Equation 6 is a cross-entropy loss applied locally among part of classes. To be noticed, MAE loss tends to oscillate near the optimal point while cross-entropy loss can always be minimized by larger class weight norms (e.g., by scaling). Thus, it is natural to find the norms of new class weights with LCE loss are bigger than those of old class weights with HKD loss in Figure 4(c).
@@
On the other hand, we attribute the bias caused by synthetic samples (as shown in Figure 4(d)) to their poor separability. Due to the limited generation quality, the synthetic sample tends to mix many features of different classes [13]. Thus it introduces noisy information when training the classifier $g(\cdot)$ and resultantly causes a large variance in the gradient direction of $g(\cdot)$ with different samples. Assume there are $M$ batches within one epoch. At each batch $B^{m}$, the derivative of $\mathcal{L}_{lce}^{m}$ with respect to $\boldsymbol{w_{k}}$ is $\frac{\partial\mathcal{L}_{lce}^{m}}{\partial\boldsymbol{w_{k}}}$. The gradient direction variance is reflected in the ratio of the norm of the mean gradient vector to the mean norm of gradient vectors, denoted as $R_{k}$ as follows:
@@
If the direction variance is large, $R_{k}$ tends to be small and vice versa. On top of that, we employ $\bar{R}=\frac{1}{|C|}\sum_{k\in C}R_{k}$ as the metric for measuring the variance regarding the norms of all class vectors. Notably, we split the dataset into batches to simulate the real scene during training. We plot $\bar{R}$ of classes belonging to synthetic and real samples in the synthetic (LCE) - real (LCE) scheme as shown in Figure 5. It shows that the gradient direction variance of class weights of real data is smaller than that of synthetic data in the whole training stage. It is conducive to a fast convergence and stable promotion of the weight norm, which explains the phenomenon of Figure 4(d). The new observation about data quality suggests explicit debiasing strategy is preferred and debiasing methods [52, 48] focusing on split loss functions may not be compatible with synthetic data in DFCIL. Results of the comparison with different debiasing methods can be found in Section V-C.
IV-C2 Weight Alignment Regularization Loss
To correct the bias, we try to align the weights of old and new classes. Though a post-processing alignment strategy is provided in [17], it is not that suitable for data consistency enhancement and will still cause bias among tasks as shown in Section V-C. We prefer to replay data through an unbiased trained model to maximize the satisfaction of the Gaussian assumption. If we align the weights after training, the feature extractor will be biased and the tied covariance may be a biased estimation. So we regularize it during training. Denoting the norm of $||\boldsymbol{w_{k}}||_{2}$ as $n_{k}$, the mean norm of old and new class weights as $n_{old},n_{new}$, the regularization is as follows:
As a result, the overall loss for the training stage is:
V Experiments
In this section, we perform experiments to evaluate our proposed CCIL framework. First, we compare CCIL with other SOTA methods on several class incremental learning benchmarks to show its superiority. Then, we perform the ablation study on the key designs of our proposed CCIL framework and we also conduct parameter analysis experiments.
V-A Experimental Settings
TABLE II: Performance on CIFAR-100. The dataset is divided into 5, 10 and 20 tasks. * indicates the results reported in [13].
Method 5 10 20
$A_{N}$ $\bar{A}$ $A_{N}$ $\bar{A}$ $A_{N}$ $\bar{A}$
DGR* 14.40 $\pm$ 0.40 - 8.10 $\pm$ 0.10 - 4.10 $\pm$ 0.30 -
LwF* 17.00 $\pm$ 0.10 - 9.20 $\pm$ 0.00 - 4.70 $\pm$ 0.10 -
DeepInversion* 18.80 $\pm$ 0.30 - 10.90 $\pm$ 0.60 - 5.70 $\pm$ 0.30 -
CCIL(ABD) 50.27 $\pm$ 0.51 64.89 $\pm$ 1.24 40.01 $\pm$ 0.86 58.86 $\pm$ 2.00 25.46 $\pm$ 0.40 46.85 $\pm$ 1.43
CCIL(R-DFCIL) 52.46 $\pm$ 0.35 66.31 $\pm$ 1.46 43.69 $\pm$ 0.57 60.14 $\pm$ 1.87 31.72 $\pm$ 0.22 49.01 $\pm$ 1.81
TABLE III: Performance on Tiny-ImageNet. The dataset is divided into 5, 10 and 20 tasks. * indicates the results reported in [13].
Method 5 10 20
$A_{N}$ $\bar{A}$ $A_{N}$ $\bar{A}$ $A_{N}$ $\bar{A}$
CCIL(ABD) 33.04 $\pm$ 0.30 46.95 $\pm$ 0.73 25.45 $\pm$ 0.48 42.56 $\pm$ 1.05 16.92 $\pm$ 0.56 36.09 $\pm$ 0.63
CCIL(R-DFCIL) 36.89 $\pm$ 0.73 49.69 $\pm$ 0.76 30.90 $\pm$ 0.52 45.29 $\pm$ 0.73 24.54 $\pm$ 0.27 39.65 $\pm$ 0.19
TABLE IV: Performance on ImageNet-100. The dataset is divided into 5, 10 and 20 tasks. * indicates the results reported in [13].
Method 5 10 20
$A_{N}$ $\bar{A}$ $A_{N}$ $\bar{A}$ $A_{N}$ $\bar{A}$
ABD 52.04 67.00 38.34 58.08 21.74 44.69
R-DFCIL* 53.10 68.15 42.28 59.10 30.28 47.33
R-DFCIL 50.90 67.72 41.38 58.82 27.86 45.74
CCIL(ABD) 56.02 69.42 41.86 60.52 25.88 48.10
CCIL(R-DFCIL) 56.80 70.39 46.04 62.66 34.60 49.18
$A_{i}=\frac{1}{|X_{test}^{0:i-1}|}\sum_{(x,y)\in X_{test}^{0:i-1}}\mathbb{I}(\mathop{\mathrm{argmax}}\limits_{k}f_{0:i-1}(x)_{k}=y),$ (17)
where $\mathbb{I}(\cdot)$ is the indicator function.
The basic training settings of CCIL are kept the same as baselines for fair comparison [12, 13]. Specifically, the backbone network for CIFAR-100 and Tiny-ImageNet is a 32-layer ResNet [3]. And for ImageNet-100, we adopt Resnet18 [3] as the backbone. At each task of CIFAR-100 or Tiny-ImageNet, we train the model for 200 epochs. The learning rate is set at 0.1 initially and decayed by 10 at epoch 80 and 120. The weight decay is 0.0005 for CIFAR-100 and 0.0002 for Tiny-ImageNet. We use SGD optimizer and the batch size is set as 128. For ImageNet-100, we train the model for 120 epochs. The learning rate is 0.1 and decayed at epoch 30 and 60 and the weight decay is set as 0.0001. We use SGD optimizer and the batch size is set as 64. For the generation, we use different generative models for different datasets to generate images of the same size as the real ones. We use the Adam optimizer with a constant learning rate 0.001. We train the generative model for 5000 steps when synthesizing images of CIFAR and Tiny-ImageNet and we train for 10000 steps when synthesizing images of ImageNet.
V-B Incremental Learning Performance
To validate the effectiveness of the proposed CCIL method, we compare it with several baselines. We briefly introduce them as follows:
Always Be Dreaming [12] (ABD): a strong baseline, which keeps the inversion stage almost the same as [14] and changes the training stage for better performance.
Because CCIL focuses on the inversion stage and the regularization of the classifier, which is orthogonal to SOTA works, we combine it with them (R-DFCIL and ABD) for evaluation. To make a fair comparison, we keep the specific training settings of CCIL the same as the combined baseline. If not explicitly mentioned, CCIL refers to CCIL + R-DFCIL in the following. We report the results of CIFAR-100 in Table II, the results of Tiny-ImageNet in Table III, and the results of ImageNet-100 in Table IV. From all three tables, we can see that our CCIL method can achieve consistent improvements both in $A_{N}$ and $\bar{A}$ compared to the baselines. Specifically, CCIL surpasses the second-best method by 2.5, 1.1, and 1.4 percent in $A_{N}$ on CIFAR-100. Likely, the absolute gains are 1.6, 0.9, and 0.5 on Tiny-ImageNet, and the improvements achieve 5.9, 4.7, and 6.7 on ImageNet-100. What is more, we can see that both R-DFCIL and ABD can benefit from the CCIL framework consistently, which verifies the applicability of CCIL. In addition, the results in $\bar{A}$ are consistent with results in $A_{N}$, which indicates that CCIL can improve performance in the whole learning process.
TABLE V: Comparisons with other debiasing methods on CIFAR-100. All numbers in the table are last incremental accuracy (mean±$\pm$±std%).
Method 5 10 20
SCE 47.86 $\pm$ 0.51 36.00 $\pm$ 0.55 22.43 $\pm$ 0.76
ACE 47.11 $\pm$ 0.32 35.13 $\pm$ 0.91 21.75 $\pm$ 0.83
SSIL 51.52 $\pm$ 0.31 42.81 $\pm$ 0.90 28.96 $\pm$ 0.11
WA 51.08 $\pm$ 0.58 42.41 $\pm$ 0.59 29.81 $\pm$ 0.32
CCIL 52.46 $\pm$ 0.35 43.69 $\pm$ 0.57 31.72 $\pm$ 0.22
V-C Comparisons with Other Debiasing Approaches
In Section IV-C, we identify two reasons for class weight bias encountered in the inversion-based incremental learning process and propose a simple regularization term $L_{war}$ to align the weights. In this section, we compare the proposed strategy with some known approaches in the literature that can help reduce bias. We will briefly describe these approaches as follows:
The approaches above can be divided into two classes: (1) split losses for data of old and new classes or even for data of each task (SCE, ACE, and SSIL), (2) explicit post-processing class weight alignment (WA). For the first line of approaches, we keep the inversion stage the same as CCIL and replace $L_{hkd}$ with the proposed losses for fair comparisons. For WA, we keep both the inversion and the training stage the same as CCIL and align the weight after the training stage.
Figure 6: The norms of class weights in 5-task experiments with different debiasing approaches. (a) Asymmetric Cross-entropy. (b) Separated-Softmax for Incremental Learning. (c) Weight Aligning. (d) CCIL with Weight Alignment Regularization.
We compare these methods by both the weight alignment situation and the final classification performance. The experiments are carried out on the CIFAR-100 dataset and the classification performance is reported in Table V. In addition, we observe the weight alignment situation at the end of the last task in the 5-task setting, as shown in Figure 6. Note that SCE performs similarly to ACE and the 2-task situation of SCE can be obtained by referring to Figure 4(d).
From the results, we can see that SCE and ACE can mitigate the weight bias among old classes but can not mitigate the bias between new and old classes due to the data reason detailed in Section IV-C. What is more, lacking information on non-target classes is attributed to the large drop in performance. SSIL intends to separate classes by tasks and keep information on non-target classes when distillation. Thus, though existing bias like SCE, it would not affect the performance much. As for WA, it performs the debiasing between old and new classes but lacks constraint within old classes, where the bias can be observed. What is more, the way of post-processing can not guarantee debiasing in the training stage which can result in a biased feature extractor and hurt the data consistency enhancement by contradicting the assumption. These factors incur the suboptimal performance of WA. Different from them, our CCIL avoids these disadvantages and mitigates the bias explicitly. However, CCIL does not keep the relative magnitude of the norms of the weight vectors within the task. We guess that task-wise alignment rather than class-wise alignment in CCIL is enough to mitigate the bias, which we will leave for future work.
V-D Ablation Study
TABLE VI: Ablation Study. We ablate each component of CCIL. All numbers in the table are last incremental accuracy (mean±$\pm$±std%).
Method 5 10 20
R-DFCIL 49.90 $\pm$ 0.23 42.57 $\pm$ 0.71 30.35 $\pm$ 0.12
CCIL--DCE 50.95 $\pm$ 0.18 42.96 $\pm$ 0.98 31.08 $\pm$ 0.32
CCIL--WAR 50.86 $\pm$ 0.34 43.52 $\pm$ 0.99 31.27 $\pm$ 0.19
CCIL 52.46 $\pm$ 0.35 43.69 $\pm$ 0.57 31.72 $\pm$ 0.22
The two main contributions of CCIL are the proposed data consistency-enhanced (DCE) loss and the weight alignment regularization (WAR) loss. The DCE loss is applied in the inversion stage and the WAR loss is used in the training stage. To validate the effectiveness of every component of the proposed CCIL, we ablate DCE and WAR respectively. As a result, we have 4 combinations: (1) CCIL. The complete framework with both loss terms, (2) CCIL--WAR. CCIL framework without the WAR loss in the training stage, (3) CCIL--DCE. CCIL framework without DCE loss in the inversion stage, (4) R-DFCIL. CCIL degrades to R-DFCIL without both DCE and WAR loss. We conduct experiments on CIFAR-100 as shown in Table VI.
V-E Parameter Analysis
In the CCIL framework, we keep the hyperparameters the same as them in [13] and there left two new hyperparameters $\lambda_{dce}$ and $\lambda_{war}$ to tune. We conduct the parameter analysis on CIFAR-100 with 5, 10, and 20 tasks. Because $\lambda_{dce}$ and $\lambda_{war}$ are introduced in the different stages, we perform two separate searches. First, we fix $\lambda_{war}=0.1$ and search the optimal $\lambda_{dce}$. We tried $\lambda_{dce}=0.01,0.02,0.05,0.1,0.2$. As shown in the left subfigure of Figure 7, the last incremental accuracy first increases and then decreases as $\lambda_{dce}$ increases, which indicates that a local optimum exists. To be noticed, CCIL performs better than the corresponding baseline in most cases. Then we fixed $\lambda_{dce}=0.05$ and search $\lambda_{war}$ in the range of $\{0.02,0.05,0.1,0.2,0.5\}$. Results in Figure 7 demonstrate that the performance is relatively less sensitive to $\lambda_{war}$ than $\lambda_{dce}$. The performance gain is close except when the regulation strength is set too high and the number of tasks is 20. Similarly, a wide range of values of $\lambda_{war}$ has positive effects on the final performance. According to the results, we set $\lambda_{dce}=0.05,\lambda_{war}=0.1$ to fit CIFAR-100 with different numbers of tasks and further adopt them in Tiny-ImageNet and ImageNet-100 experiments. More surprisingly, we also adopt the same parameters in the CCIL + ABD experiments and it works well. Results in Section V-B validate the good transferability of the parameters.
VI Conclusion

### v2

@@ 5.6 Quantitative Measurement of Data Consistency
Table 6: Effects of Different Losses on Data Consistency. We ablate loss components of the baseline loss and show the improvement after combining $L_{dce}$.
@@
In this section, we give the quantitative measurement of data consistency with respect to different loss combinations. We adopt both parametric estimation and non-parametric estimation to estimate $p(\boldsymbol{z})$ and $q(\boldsymbol{z})$ for a comprehensive understanding. For parametric estimation, we use the modeling in Section 4.2.1, while for non-parametric estimation, we employ the kernel density estimation (KDE) technique. Specifically, we use the Gaussian kernel and decide the bandwidth by Scott’s method [56]. With estimated $p(\boldsymbol{z})$ and $q(\boldsymbol{z})$, one can calculate $D_{KL}$. But owing to no analytical solution and high computational complexity, we approximate KL divergence with Monte-Carlo approximation [^1] We follow the algorithm in http://joschu.net/blog/kl-approx.html.. With the two KL Divergence metrics, we try to study the effects of loss combinations proposed in the literature under a unified framework. We first train a model with 50 classes in CIFAR-100 and then invert samples from it. Then, the KL divergences of inverted and real old data approximated in two ways are given. We ablate the loss components and the results can be found in Table 6.
@@
We can draw some conclusions from Table 6. First, the results of $D_{KL}$ (KDE) and $D_{KL}$ (Gaussian) are similar, which corroborate each other and can reflect the relative data consistency among methods. Second, the cross-entropy loss does not contribute to data consistency much. It is intuitive because cross-entropy loss is a sample-wise measurement and lacks any direct constraint on either the diversity or the distribution. What is more, minimizing the cross-entropy loss in Equation 1 equals maximizing the Softmax score of samples. But it is well-studied that Softmax score [57] is not a good metric for OOD detection [49, 48] and thus for data consistency measurement. Note in our experiments, the cross-entropy is not indispensable. We keep it for it does not contradict other losses during training and can slightly improve the performance in some cases. Third, the statistics alignment loss indeed ameliorates the consistency, but there is still plenty of room. We can observe that the alignment in Equation 2 only considers the features of the BN layers but lacks explicit regularization on non-BN layer features. As a result, the proposed data consistency enhancement loss $L_{dce}$ on the features of the penultimate layer further enhance consistency as desired.
5.7 Analysis of Weight Bias
@@
Figure 6: The norms of class weights in 2-task experiments with different training schemes for the second task on CIFAR-100, ImageNet-100, and Tiny-ImageNet. (a) the standard training scheme. (b) a scheme where real old data is used and trained with local CE loss. (c) a scheme where the real old data is used. (d) a scheme where the synthetic old data is trained with the local CE loss. We report the results at the end of the final task.
@@
However, when we control the number of classes of synthetic and real data to be the same, the weight bias problem still exists (as shown in Figure 6(a)). It indicates the existence of other reasons. We argue different loss functions and different training data may also contribute to weight bias and thus design a two-task class incremental experiment to uncover the factors concerned with it. The first task includes data from first 50% classes while the second task includes the other 50% classes. We conduct experiments on CIFAR-100, ImageNet-100, and Tiny-ImageNet. Training with standard CE is adopted for the first task. As for the second task, we compare 4 training schemes: (1) synthetic (HKD) - real (LCE), which is the standard training scheme, (2) real (LCE) - real (LCE), where real old data is used and trained with local CE loss, (3) real (HKD) - real (LCE), where the real old data is used, (4) synthetic (LCE) - real (LCE), where the synthetic old data is trained with the local CE loss. We report the norm of class weights at the end of the second task.
First of all, Figure 6(b) shows an unbiased situation under the condition of equal losses and data quality for both tasks. Further, results in Figure 6(c) and (d) validate our hypothesis that both the different training losses and different properties of training data will have effects on the class weights. HKD loss calculates the mean square error (MAE) of the logits of the old and new models as defined in Equation 5. LCE loss defined in Equation 6 is a cross-entropy loss applied locally among part of classes. To be noticed, MAE loss tends to oscillate near the optimal point while cross-entropy loss can always be minimized by larger class weight norms (e.g., by scaling). Thus, it is natural to find the norms of new class weights with LCE loss are bigger than those of old class weights with HKD loss in Figure 6(c) on CIFAR-100 and ImageNet-100 datasets. However, the situation changes on the Tiny-ImageNet dataset. We attribute it to the low classification accuracy, which indicates that the gradient direction of class weight vectors is not consistent. As a result, the norm will not increase consistently. In contrast, the gradient of HKD loss is more stable, resulting in larger norms.
@@
On the other hand, we attribute the bias caused by synthetic samples (as shown in Figure 6(d)) to their poor separability. Due to the limited generation quality, the synthetic sample tends to mix many features of different classes [15]. Thus it introduces noisy information when training the classifier $g(\cdot)$ and resultantly causes a large variance in the gradient direction of $g(\cdot)$ with different samples. Assume that there are $M$ batches within one epoch. At each batch $B^{m}$, the derivative of $\mathcal{L}_{lce}^{m}$ with respect to $\boldsymbol{w_{k}}$ is $\frac{\partial\mathcal{L}_{lce}^{m}}{\partial\boldsymbol{w_{k}}}$. The gradient direction variance is reflected in the ratio of the norm of the mean gradient vector to the mean norm of gradient vectors, denoted as $R_{k}$ as follows:
@@
If the direction variance is large, $R_{k}$ tends to be small and vice versa. On top of that, we employ $\bar{R}=\frac{1}{|C|}\sum_{k\in C}R_{k}$ as the metric for measuring the variance regarding the norms of all class vectors. Notably, we split the dataset into batches to simulate the real scene during training. We plot $\bar{R}$ of classes belonging to synthetic and real samples in the synthetic (LCE) - real (LCE) scheme as shown in Figure 7. It shows that the gradient direction variance of class weights of real data is smaller than that of synthetic data in the whole training stage. It is conducive to a fast convergence and stable promotion of the weight norm, which explains the phenomenon of Figure 6(d). The new observation about data quality suggests explicit debiasing strategy is preferred and debiasing methods [55, 52] focusing on split loss functions may not be compatible with synthetic data in DFCIL. Results of the comparison with different debiasing methods can be found in Section 5.3.
5.8 Computational Overhead
Due to the existence of the estimation stage, it needs more computation than the baseline method. In this section, we report the exact computation overhead. Denote the forward-pass overhead of a data batch through the classification model as $c_{p}$, the backward-pass overhead as $c_{b}$, the forward-pass overhead of a noise batch through the generator as $g_{p}$, the backward-pass overhead as $g_{b}$, the number of iterations of the inversion stage is $N_{inv}$, the number of epochs in the training stage is $N_{e}$, and the number of batches in one epoch is $M$. Then, we can obtain the computational overhead for the inversion stage, training stage, and estimation stage as $N_{inv}\times(g_{p}+g_{b}+c_{p}+c_{b})$, $N_{e}\times M\times(g_{p}+2\times(c_{p}+c_{b}))$, and $M\times(|\mathcal{C}_{old}|/|\mathcal{C}_{new}|\times(g_{p}+c_{p})+c_{p})$. In the settings of our experiments, $|\mathcal{C}_{old}|/|\mathcal{C}_{new}|\ll N_{e}$ and $M\times|\mathcal{C}_{old}|/|\mathcal{C}_{new}|\ll N_{inv}$. As a result, the extra computational overhead of the estimation stage is negligible. We show the practical running time in Table 7, which is consistent with the theoretical complexity.
Table 7: Computational overhead on different benchmarks. $V_{a}^{b}$ denotes the cost for ER with $a$ tasks on $b$ dataset.
CIFAR-100 Tiny-ImageNet ImageNet-100
5 10 20 5 10 20 5 10 20
ER $V_{5}^{C}$ $V_{10}^{C}$ $V_{20}^{C}$ $V_{5}^{T}$ $V_{10}^{T}$ $V_{20}^{T}$ $V_{5}^{I}$ $V_{10}^{I}$ $V_{20}^{I}$
[…]
6 Limitations
One limitation of CwD is the increased computational cost compared to non-data-free baselines like ER. The introduction of both the inversion stage and the estimation stage in each task results in higher computational demands. Detailed information on the computational overhead is provided in Section 5.8. Another potential limitation is the restricted modeling of the data distribution. Currently, our modeling approach is limited to the features of the penultimate layer under a multivariate Gaussian assumption. However, statistics from other layers may also contribute to addressing the inversion challenge. Exploring and incorporating these additional statistics is a promising direction for future research and may lead to improved performance. Lastly, by applying $\mathcal{L}_{war}$ in the training stage, CwD does not preserve the relative magnitudes of weight vector norms within the task. We hypothesize that task-wise alignment, rather than class-wise alignment, may suffice to mitigate bias in CwD, which we plan to investigate in future work.
7 Conclusion